31. August 2015: Monday

LDAP/IdM:

  • LDAP Updates:
    • Several issues Friday night, dealt with them
    • One of them though is unclear – asking payroll to confirm.
    • Merged two accounts per request.
  • ACCOUNT Updates:
    • No issues.
  • VPN Updates:
    • Have a concern that the EAS controlled VPN memberships are in danger of being clobbered, but I’m not certain of that.  Need to test that possibility out.

Backups:

  • Save Groups:
    • Failed: windowpane (are you surprised???)
  • Clones:
    • waiting for tape relabel to finish to prevent reuse of tapes I don’t quite want reused yet.
  • Problems:
    • A0945SG: parts were delivered to the mail room – I tracked them down and retrieved them late Friday afternoon.  Parts installed – and the new parts work!  YAY.  And then I rebooted and nothing works 🙁 And then we found out the brocade config was messed up – and we got it working better than ever 😉
  • Tapes:
    • Getting those 168 tapes schlepped back from the Vault on Friday deposited and relabeled.
      • deposit accomplished.
      • relabel started.

Calendar:

  • Hush Hush Audit Stuff.
  • Digging into a problem with a Calendar Resrource – no idea what has gone wrong.

Tags:

Leave a Reply

You must be logged in to post a comment.